OpenStudy (anonymous):

Equation ... Pls solve . N explain. x(4x-3)=(2x-1)^2.

OpenStudy (anonymous):

x(4x-3) = (2x-1)^2 => 4x^2-3x = 4x^2-4x+1 => 4x^2-4x^2 =3x-4x+1 => 0 = -x+1 => x=1

OpenStudy (anonymous):

Distribute the x inside the (). 4x^2-3x Square (2x-1) using FOIL. 4x^2-6x+1 4x^2-3x = 4x^2 - 6x +1 4x^2 cancels from both sides -3x = -6x + 1 3x = 1 x = 1/3

OpenStudy (anonymous):

your foil should be 4x^2-4x+1

OpenStudy (anonymous):

You're right. I should stop I've been sitting here too long

OpenStudy (anonymous):

haha no you're fine